OpenStudy (anonymous):

the set of ordered pairs represent a function, write a rule that represents the function. (0,0) (1,4) (2,16) (3,36) (4,64) y= simplify type an expression using x as the variable

OpenStudy (anonymous):

nope (0,0)

OpenStudy (anonymous):

ok then try \(y=4x^2\)

OpenStudy (anonymous):

since \(4\times 0^2=0,4\times 1^2=4, 4\times 2^2=16...\)

OpenStudy (anonymous):

thanks